A ball is thrown upward with a velocity of 29.4m/search. From a tower that is 98m/s: calculate 1.time it takes to reached the height point. 2. The height distance reached. 3. Time it takes to reached the ground. ( g-9.8m/s)

Grade:7

1 Answers

Rishav Ranjan
6 years ago
You can do it by a simple formulau=29m/s and g=9.8 m/s^21) now time to reach maximum height=u/g=29.4/9.8=3s.2) maximum height=u^2/g=29.4×29.4/2×9.8=44.1m3)total time=2u/g=2×29.4/9.8=6s.